# High Resolution Millimeter Wave Imaging Based on FMCW Radar Systems at W-Band

### v1

@@ Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, low-weight, and highly accurate millimeter wave (mm-Wave) imagery system capable of operating in all weather conditions. We describe mm-Wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this work, can strongly compete with Lidar systems as the reolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
@@
FMCW radar, high resolution imaging, mm-Wave imaging, SAR.
@@ I Introduction
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 18, 5, 6, 7, 8, 9, 10]. Compared to the systems operating at lower frequencies, the advantage of the systems operating at mm-Wave, sub-terahertz, and terahertz frequencies is that they are light-weight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. Also, operating at higher frequencies makes it possible to create apertures with large number of elements while they occupy small physical space. This, in turn, yields high resolution in the azimuth direction.
The imagery systems operating at millimeter wave (mm-Wave), sub-terahertz, and terahertz frequencies have numerous applications including material characterization, high resolution localization, non-destructive testing (NDT), and near-field imaging, to mention a few.
@@ III Apodization
The ADC block contains the ADC as well as the low pass Anti Aliasing Filter (AAF).
The sampling frequency for the ADC is 5 MHz. The chirp time has been set to 5 msec.
@@
Figure 7: The radar system along with a two-axis motorized scanner to create the synthetic aperture.
Fig. 7 illustrates the radar system along with a two-axis motorized scanner to create the synthetic aperture.
@@ VII-F More Results
Figure 21: a) the experimental set-up presented in Fig. 20-(a), the data has been collected in a foggy condition, b) the reconstructed image.
@@ VIII Conclusion
In this paper, we presented our custom-made FMCW radar system which operates at mm-Wave frequencies. We described the functionalities of the system in detail. We further, presented the imaging algorithm and several signal processing techniques for image formation and image post processing. At the end, we presented all the results based on the experimental data gathered from the imagery system. The paper is a proof that images with high resolution and high quality can be generated using compact, low-weight, and low-cost mm-Wave systems such as the one presented in this work.

### v2

@@ Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, low-weight, and highly accurate millimeter wave imagery system capable of operating in all weather conditions. We describe millimeter wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this work, can strongly compete with Lidar systems as the resolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
@@
FMCW radar, high resolution radar imaging, SAR.
@@ I Introduction
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 18, 5, 6, 7, 8, 9, 10]. Compared to the systems operating at lower frequencies, the advantage of the systems operating at millimeter wave, sub-terahertz, and terahertz frequencies is that they are light-weight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. Also, operating at higher frequencies makes it possible to create apertures with large number of elements while they occupy small physical space. This, in turn, yields high resolution in the azimuth direction.
The imagery systems operating at millimeter wave, sub-terahertz, and terahertz frequencies have numerous applications including material characterization, high resolution localization, non-destructive testing (NDT), and near-field imaging, to mention a few.
@@ III Apodization
The ADC block contains the ADC as well as the low pass Anti Aliasing Filter (AAF). The sampling frequency for the ADC is 5 MHz. The chirp time has been set to 5 msec.
@@
Figure 7: The radar system as well as the motorized scanner to create the synthetic aperture.
Fig. 7 illustrates the radar system as well as the motorized scanner to create the synthetic aperture.
@@ VII-F More Results
Figure 21: a) the experimental set-up presented in Fig. 20-(a), the data has been collected in dense fog, b) the reconstructed image.
@@ VIII Conclusion
In this paper, we presented our custom-made FMCW radar system which operates at millimeter wave frequencies. We described the functionalities of the system in detail. We further, presented the imaging algorithm and several signal processing techniques for image formation and image post processing. Finally, we presented all the results based on the experimental data gathered from the imagery system. The paper is a proof that images with high resolution and high quality can be generated using compact, low-weight, and low-cost millimeter wave systems such as the one presented in this work.